The sum of the angle measures of any triangle is 180°.find each of the angle measures of a triangle if the second angle measures 10°more than twice the first,and the third angle measures 10°moree than the second.
Answer by Alan3354(69443)

find each of the angle measures of a triangle if the second angle measures 10°more than twice the first,and the third angle measures 10°moree than the second.
-----------
A = the 1st angle
2A + 10 = 2nd angle
2A + 20 = 3rd angle
------------------------ Add
5A + 30 = 180
etc